OpenStudy (kidthatbro8):

Mike is working on solving the exponential equation 37^x = 12; however, he is not quite sure where to start. Using complete sentences, describe to Mike how to solve this equation.

OpenStudy (agent0smith):

Any idea what the first step might be?

OpenStudy (kidthatbro8):

i thought about creating equal bases...

OpenStudy (kidthatbro8):

@agent0smith ??

OpenStudy (agent0smith):

Good try... can you get 37 and 12 to be equal bases? You might notice that 37 is prime...

OpenStudy (kidthatbro8):

yeah. what do we do then?

OpenStudy (agent0smith):

You must have learned some other methods to solve these...

OpenStudy (wolf1728):

How about taking logarithms of both sides?

OpenStudy (kidthatbro8):

like, making them both into logarithms?

OpenStudy (wolf1728):

Yes, do you know how to do that?

OpenStudy (kidthatbro8):

yes,, the formula is y = b^x <---> x = logb y

OpenStudy (kidthatbro8):

so would it be something like x = log37 12? or is that wrong?

OpenStudy (wolf1728):

Here's how I'd do it 37^x = 12 Taking logs of both sides x * log (37) = log (12) x = log (12) / log (37) I think you can figure it out from there.

OpenStudy (kidthatbro8):

uh, i don't think i know how to work it out from there, could you help me?

OpenStudy (wolf1728):

Sure Log (37) = 1.5682017241 Log (12) = 1.079181246 So x = 1.079181246 / 1.5682017241 x = 0.6881648129 To check that out do you have a calculator that can calculate 37^0.6881648129 ?

OpenStudy (kidthatbro8):

my calculator says it's 11.9999999996

OpenStudy (wolf1728):

Pretty darned close to 12 isn't it ? :-)

OpenStudy (kidthatbro8):

oh, yeah!

OpenStudy (wolf1728):

So you see x = 0.6881648129

OpenStudy (kidthatbro8):

alright, thank you so much!!

OpenStudy (wolf1728):

You are welcome. I know taking logs of both sides seems pretty weird at first - but that's how you get the answer!